OpenStudy (anonymous):

.Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3).

OpenStudy (anonymous):

y=2x+2 given ......so for perpendicular i.e.....-1/2....m=-1/2 y-3=-1/2(x-6)....solving it...2y-6=-x+6.......so....x+2y-12=0....